OpenStudy (anonymous):

Which is an equation in slope-intercept form for the line passing through (–1, 5) and (5, 8)

OpenStudy (the_fizicx99):

Find the slope. M = (y2-y1) / (x2-x1) then input your values into: y-y1=m(x-x1)